# MangaVQA and MangaLMM: A Benchmark and Specialized Model for Multimodal Manga Understanding

### v2

@@ Abstract
Jeonghun Baek††thanks: Equal contribution. Kazuki Egashira11footnotemark: 1 Shota Onohara11footnotemark: 1 Atsuyuki Miyai11footnotemark: 1 Yuki Imajuku Hikaru Ikuta Kiyoharu Aizawa The University of Tokyo baek@hal.t.u-tokyo.ac.jp https://github.com/manga109/MangaLMM/
@@ 3.2 MangaOCR: A Consolidated Dataset for Manga Text Recognition
Count type Total Train Valid Test Comic volumes 109 89 7 13 Images 10,602 8,763 673 1,166 MangaOCR Dialogue 148K 120K 9K 18K Onomatopoeia 61K 50K 4K 7K Total 209K 170K 13K 26K MangaVQA QA pairs 40,363 39,837 526
@@ 5.1 Training Data Construction
Synthetic VQA Training Set . For the VQA task, we generate synthetic training data using GPT-4o gpt4o (gpt-4o-2024-11-20). Following the synthetic data construction used in LLaVA llava, we generate five questions per image using both the image and its annotation from the OCR training set . Here we exclude < 0.1% of the images where the text annotation is not included or GPT-4o refused to respond (e.g., due to violent content). Although we requested GPT-4o to generate five questions per image, it occasionally returned fewer than five. As a result, we created a total of 39,837 synthetic VQA samples from 8,379 images. The prompt used for question generation is provided in the supplementary materials. We plan to release this as a training split of our MangaVQA.
@@ 5.2 Training Details
Training Strategy. We perform continual finetuning on both and using the pretrained Qwen2.5-VL 7B (Qwen2.5-VL-7B-Instruct). Most hyperparameters follow the original Qwen2.5-VL configuration, with a few modifications. For Manga109 images (16541170 resolution), we follow Qwen2.5-VL’s image resizing mechanism, which is based on pixel count thresholds, where the minimum and maximum number of input pixels are 3,136 and 2,116,800, respectively.
@@ 6 Experiments
Evaluation Protocol for MangaVQA. Following LLaVA-Bench llava, we adopt the LLM-as-a-judge approach judge as our evaluation metric. We select a single LLM as the judge, which assigns scores to model responses. Specifically, we provide the judge with the question, a human-written answer, and the model’s response. Based on the human-written answer, the judge evaluates whether the model’s response is appropriate and relevant to the question on a 1–10 scale. To avoid circular bias with GPT-4o, which was used in constructing our VQA training dataset, we employ Gemini 2.5 Flash gemini2.5 (gemini-2.5-flash) as the judge LLM. The prompt used for LLM-as-a-judge is provided in the supplementary materials.
@@ 6.1 Main Results
MangaOCR MangaVQA FT data Hmean (%) LLM (/10.0) None 0.9 5.65 74.5 1.3 1.20 0.32 0.0 0.0 6.57 0.06 + 71.2 0.6 6.68 0.03
@@ Acknowledgments
In this supplementary material, we provide additional details including (A) OCR evaluation in comics, (B) synthetic VQA examples, (C) additional MangaVQA examples, (D) setup details, and (E) additional results.
@@ Appendix B Synthetic VQA Examples
Human Validation. To validate the reliability of the synthetic VQA data generated by GPT, we conducted a manual evaluation. We randomly sampled 100 question–answer pairs and asked four human evaluators to assign scores to each item on a three-level scale: 0 (incorrect), 0.5 (partially correct), and 1 (correct). The average score is 0.78, suggesting that approximately 80% of the synthetic VQAs are judged to be appropriate by humans.
@@ E.2 More Analysis of MangaVQA
To validate the reliability and consistency of the Gemini-judge employed in the MangaVQA evaluation, we conducted a comparative analysis between its evaluation scores and those provided by human annotators. Specifically, we randomly sampled 100 items from the benchmark dataset and asked two human evaluators to assign scores to each item, following the same evaluation prompt used for the Gemini-judge.

### v3

@@ Abstract
Jeonghun Baek††thanks: Equal contribution. Kazuki Egashira11footnotemark: 1 Shota Onohara11footnotemark: 1 Atsuyuki Miyai11footnotemark: 1 Yuki Imajuku Hikaru Ikuta Kiyoharu Aizawa The University of Tokyo baek@hal.t.u-tokyo.ac.jp https://manga109.github.io/MangaVQA_LMM/
@@ 3.2 MangaOCR: A Consolidated Dataset for Manga Text Recognition
| Count type | Total | Train | Valid | Test |
| Comic volumes | 109 | 89 | 7 | 13 |
| Images | 10,602 | 8,763 | 673 | 1,166 |
| MangaOCR | ||||
| Dialogue | 148K | 120K | 9K | 18K |
| Onomatopoeia | 61K | 50K | 4K | 7K |
| Total | 209K | 170K | 13K | 26K |
| MangaVQA | ||||
| QA pairs | 40,363 | 39,837 | 526 |
@@ 5.1 Training Data Construction
Synthetic VQA Training Set . For the VQA task, we generate synthetic training data using GPT-4o gpt4o (gpt-4o-2024-11-20). Following the synthetic data construction used in LLaVA llava, we generate five questions per image using both the image and its annotation from the OCR training set . Here we exclude < 0.1% of the images where the text annotation is not included or GPT-4o refused to respond (e.g., due to violent content). Although we requested GPT-4o to generate five questions per image, it occasionally returned fewer than five. As a result, we created a total of 39,837 synthetic VQA samples from 8,379 images. The prompt used for question generation is shown in Table A in the Appendix. We plan to release this as a training split of our MangaVQA.
@@ 5.2 Training Details
Training Strategy. We perform continual finetuning on both and using the pretrained Qwen2.5-VL 7B (Qwen2.5-VL-7B-Instruct). Most hyperparameters follow the original Qwen2.5-VL configuration, with a few modifications. For Manga109 images (16541170 resolution), we follow Qwen2.5-VL’s image resizing mechanism, which is based on pixel count thresholds, where the minimum and maximum number of input pixels are 3,136 and 2,116,800, respectively. We train for one epoch with a batch size of 32.
@@ 6 Experiments
Evaluation Protocol for MangaVQA. Following LLaVA-Bench llava, we adopt the LLM-as-a-judge approach judge as our evaluation metric. We select a single LLM as the judge, which assigns scores to model responses. Specifically, we provide the judge with the question, a human-written answer, and the model’s response. Based on the human-written answer, the judge evaluates whether the model’s response is appropriate and relevant to the question on a 1–10 scale. To avoid circular bias with GPT-4o, which was used in constructing our VQA training dataset, we employ Gemini 2.5 Flash gemini2.5 (gemini-2.5-flash) as the judge LLM. The prompt used for LLM-as-a-judge is shown in Table C in the Appendix.
@@ 6.1 Main Results
| MangaOCR | MangaVQA | |
| FT data | Hmean (%) | LLM (/10.0) |
| None | 0.9 | 5.65 |
| 74.5 1.3 | 1.20 0.32 | |
| 0.0 0.0 | 6.57 0.06 | |
| + | 71.2 0.6 | 6.68 0.03 |
@@ Appendix B Synthetic VQA Examples
Human Validation. To validate the reliability of the synthetic VQA data generated by GPT, we conducted a manual evaluation. We randomly sampled 500 question–answer pairs and asked four human evaluators to assign scores to each item on a three-level scale: 0 (incorrect), 0.5 (partially correct), and 1 (correct). The average score is 0.80, suggesting that approximately 80% of the synthetic VQAs are judged to be appropriate by humans.
@@ E.2 More Analysis of MangaVQA
To validate the reliability and consistency of the Gemini-judge employed in the MangaVQA evaluation, we conducted a comparative analysis between its evaluation scores and those provided by human annotators. Specifically, we asked two human evaluators to assign scores to all items in the benchmark dataset, following the same evaluation prompt used for the Gemini-judge.
@@ E.3 More Analysis of MangaOCR
Comparison with dedicated OCR system.. We evaluated DeepSolo++ ye2023deepsolo++, a multilingual OCR system, on the MangaOCR task. The model achieved an Hmean of 5.4%, indicating very low performance in this domain. Since DeepSolo++ operates zero-shot on manga-style images, we observed that it often detects only small fragments of the text (e.g., one or two characters) rather than identifying the full text in the speech balloons, leading to many incorrect predictions. This behavior is consistent with what we observe in other LMMs when applied zero-shot to manga.
@@ E.4 Exploring the Capability for Manga and Natural Image Understanding
Method Total (100 images) JPN subset (6 images) MangaOCR Qwen2.5-VL 7B 2.8 0.3 0.9 MangaLMM 13.5 50.0 71.5
Evaluation on Other Comics. We find that MangaLMM shows some degree of generalization to other comics, although its performance is naturally lower than on in-domain Japanese manga. As discussed in the related work, there is currently no established VQA dataset for the comic domain, making it difficult to use a standard VQA benchmark in this setting. To examine cross-domain performance, we therefore conducted an additional OCR evaluation on the well-known comic dataset eBDtheque eBDtheque2013, which contains 100 images from European, American, and Japanese comics. Table K shows the results.
The performance of MangaLMM on eBDtheque is lower than on MangaOCR, reflecting the domain differences between the datasets. Notably, MangaLMM still outperforms Qwen2.5-VL, suggesting that training on manga contributes positively to OCR performance even in other comic styles. We attribute the lower performance on eBDtheque to two main factors: (1) Page format mismatch: MangaOCR consists of two-page spreads, whereas eBDtheque images are single-page. This mismatch frequently leads to shifted or misaligned predicted bounding boxes. (2) Color vs. grayscale: MangaOCR is entirely black-and-white, while more than half of eBDtheque images are in color. MangaLMM, trained on grayscale manga, shows degraded performance on colored pages. We expect both issues to be addressable by incorporating single-page layouts and color comic images during training. While MangaLMM is intentionally specialized for manga, extending it toward a general comic-capable LMM is a promising direction for future work.